Wallace Incorporated sells its products for $550 per unit. Variable costs are currently 20% of sales revenue. Fixed expenses are $198,000 per year.

What is the breakeven point in units at the current selling price?

440 units

1800 units

450 units

300 units
Answer #1

Variable expenses=550*20%=110

Contribution margin=Sales-Variable cost

=550-110=$440

Hence breakeven=Fixed cost/Contribution margin

=198,000/440

=450 units.